The quantity demAnded and the quantity supplied at R15 are 80kg and 110kg respectively. Determine the equation of demand and supply?


1
Expert's answer
2021-06-15T12:46:46-0400

Let's take equilibrium price to be 10 and quantity 95

Demand equition will take a form "P=a-bQ"

P=price

a=y intercept when p is zero

b=slope

Q=quantity demanded

"Slope=\frac{∆y}{∆x}\\\frac{15-10}{95-80}=\frac{1}{3}"

"0=a- \frac{1}{3}×80\\a=\frac {80}{30}\\p=\frac {80}{30}- \frac{1}{3}Qd"

Supply equition will take the form

"P=a+bQ"

P=price

a=x intercept when p is zero

b=slope

Q=quantity supplied

"Slope=\frac{∆y}{∆x}\\\frac{15-10}{110-95}=\frac{1}{3}"


"0=a- \frac{1}{3}×110\\a=\frac {110}{30}\\p=\frac {110}{30}- \frac{1}{3}Qs"